# LibEER: A Comprehensive Benchmark and Algorithm Library for EEG-based Emotion Recognition

### v1

@@ Abstract
EEG-based emotion recognition (EER) is garnering increasing attention due to its potential in understanding and analyzing human emotions. Recently, significant advancements have been achieved using various deep learning-based techniques to address the EER problem. However, the absence of a convincing benchmark and open-source codebase complicates fair comparisons between different models and poses reproducibility challenges for practitioners. These issues considerably impede progress in this field. In light of this, we propose a comprehensive benchmark and algorithm library (LibEER) for fair comparisons in EER by making most of the implementation details of different methods consistent and using the same single codebase in PyTorch. In response to these challenges, we propose LibEER, a comprehensive benchmark and algorithm library for fair comparisons in EER, by ensuring consistency in the implementation details of various methods and utilizing a single codebase in PyTorch. LibEER establishes a unified evaluation framework with standardized experimental settings, enabling unbiased evaluations of over ten representative deep learning-based EER models across the four most commonly used datasets. Additionally, we conduct an exhaustive and reproducible comparison of the performance and efficiency of popular models, providing valuable insights for researchers in selecting and designing EER models. We aspire for our work to not only lower the barriers for beginners entering the field of EEG-based emotion recognition but also promote the standardization of research in this domain, thereby fostering steady development. The source code is available at https://github.com/ButterSen/LibEER.
@@ I Introduction
Emotion encompasses a series of subjective cognitive experiences, serving as the primary framework for interpersonal relationships and providing the criteria for evaluating human behavior [1]. Electroencephalography (EEG) is a widely used physiological signal in emotion analysis due to its non-invasive nature, high temporal resolution, portability, and cost-effectiveness [29, 30]. Consequently, EEG-based emotion recognition (EER) has garnered increasing attention in recent years because of its robust capacity to understand emotions and has broad applicability in fields such as healthcare [31, 40], advertising [32], and education [33].
Traditional approaches primarily focus on designing or learning discriminative features, such as power spectral density (PSD) [39] and differential entropy (DE) [15], to enhance the performance of EER. Recently, with the aid of advanced neural networks, a variety of deep learning-based models [16, 18] have sprung up to recognize emotion states from EEG signals and have been one of the most promising research topics [2]. These models often aim to capture the spatiotemporal characteristics of EEG signals and can be categorized into five types: deep neural network-based (DNN), convolutional neural network-based (CNN), recurrent neural network-based (RNN), graph neural network-based (GNN), and Transformer-based methods. DNN-based methods [5, 6] are simple yet effective, employing fully connected layers and directly using the EEG signal’s features as input. CNN-based methods [8, 9, 10] treat EEG data as 2D or 3D images, where one dimension corresponds to the electrodes and the other dimension(s) represent temporal or frequency features. RNN-based methods [11], particularly the long short-term memory (LSTM) model, are suited for modeling crucial temporal information in EEG signals. GNN-based methods [12, 13, 14] are employed to model the spatial information of EEG signals across different channels. Transformer-based methods [7] are the most powerful to date, segmenting EEG signals into multiple patches and capturing dependencies between these patches using the self-attention mechanism.
Despite the significant achievements claimed by various deep learning-based models, some key issues remain underexplored in the EER research field, hindering the further development of actual progress in this area. In this paper, we identify these issues as a lack of benchmark and codebase, which makes fair comparisons difficult and leads to practitioners struggling with reproducibility.
The most critical issue pertains to the absence of a comprehensive benchmark for EER tasks, which results in non-unified, unreasonable, and ambiguous experimental settings across different models. This situation makes fair comparisons challenging and may even call into question certain conclusions. Specifically, with regard to the non-unified issue, various methods utilize differing datasets or operate under different scenarios, employ distinct pre-processing techniques (e.g., whether to conduct artifacts removal or not), feature varying ratios of training and test sets, and apply diverse data segmentation (i.e., length of samples), among other discrepancies. In terms of the unreasonable issue, certain studies employed a fixed data split strategy [15, 17, 12] -for example, utilizing the first 9 trials as training data and the final 6 trials as test data in the SEED dataset-which may result in accidental experimental outcomes. Additionally, some studies failed to satisfy cross-trial requirements in subject-dependent scenarios [34, 11], potentially leading to data leakage issues. Moreover, the majority of methods lack a validation set for model selection [13, 5], thus yielding unreliable experimental results. Regarding the ambiguous issue, many studies [7, 35, 36] do not explicitly detail how their experimental results are reported. Furthermore, data selection procedures are often unspecified, such as which session in the SEED dataset is utilized [12, 14] and how incomplete data in the HCI dataset is handled [37, 7].
An additional critical issue revolves around the absence of an open-source library for EER tasks. Despite the impressive performance demonstrated in numerous studies, the lack of publicly available code poses a significant obstacle to reproducing these outcomes [38, 5], thereby impeding researchers from building upon these studies. Furthermore, many studies fail to provide comprehensive implementation details of the models [5, 7], neglecting to include crucial information such as hyperparameters, number of network layers, activation functions, and so forth. Moreover, different studies frequently rely on distinct development platforms, such as PyTorch [19], TensorFlow [20], and MATLAB [26]. In totality, these deficiencies curtail researchers’ capacity to devise superior methods.
In this paper, we undertake the task of addressing the aforementioned critical issues and providing insights into the advancements achieved. Our approach involves the development of a comprehensive benchmark and algorithm library for EEG-based emotion recognition (referred to as LibEER). We achieve this by ensuring consistency in the implementation details of various EER methods. Firstly, we establish a fair and rational benchmark that standardizes and unifies the utilized datasets, evaluation metrics, and experimental settings. Specifically, all EER models leverage the most commonly employed datasets (SEED [22], SEED-IV [24], DEAP [21], and HCI [23]) in the field, enabling performance comparison in both subject-dependent and cross-subject scenarios. Moreover, we standardize experimental settings and methodologies, particularly by adopting a more reasonable “Train-Validation-Test” data split strategy [25, 27], widely recognized in the machine learning community for performance verification. Further, we unify the disparate length standards of EEG samples prevalent in existing research by standardizing each EEG sample length to 1s within LibEER. Additionally, we systematically implement ten representative deep learning-based EER methods with a unified training pipeline in PyTorch. This encompasses the provision of a diverse range of mainstream EEG signal preprocessing functions, evaluation protocols, data split methodologies, and experimental scenarios. To enhance the versatility of our toolkit, we furnish user-friendly interfaces for customizing experimental schemes, allowing users to easily set hyperparameters, select preprocessing methods, and modify backbone models for conducting experiments. As a result, we are able to present an accurate depiction of the current state-of-the-art results in EER.
@@
It is noteworthy that recent research [3] has made important contributions by providing a toolbox for EER tasks. However, our work is substantially different and more incremental in several key aspects. First, we systematically analyze the problems associated with fair comparison in EER tasks, a discussion notably absent in  [3]. Second, while the existing benchmark [3] includes a variety of methods, many are not specifically tailored for EER tasks. In contrast, we have selected the ten most representative deep learning EER methods based on their impact, making our benchmarks more representative of the domain. Finally, the current benchmark [3] predominantly focuses on research with open-source code. Given that many influential EER studies do not have publicly available codes, this approach has led to the exclusion of some significant models. In response, we have carefully reproduced all representative studies including those without open-source code, which was a laborious endeavor, and performed rigorous validation of the results to ensure accuracy. We posit that our benchmark and library offer novel perspectives that can enrich the field significantly.
In summary, our main contributions are the following:
@@
  • We have pioneered the establishment of the first benchmark specifically tailored for enabling fair comparisons in EEG-based emotion recognition. This benchmark not only unveils the actual progress of the EER task but also serves as a convenient reference for conducting comparative experiments.
@@
  • We have introduced LibEER as an open-source project on GitHub, flexibly customizing experimental schemes with user-friendly interfaces. It facilitates a fair and comprehensive evaluation of over ten popular models across four datasets. Functioning as both a toolbox and a platform, LibEER assists practitioners in efficiently utilizing and replicating EER methods.
@@
  • We provide comprehensive evaluations with various experimental settings, by controlling the implementation details. These experimental results offer valuable insights into the advancements made and aid researchers in the selection or design of EER models.
II Overview of EEG-based Emotion Recognition Methods
In this section, we begin by presenting the problem formulation of EEG-based emotion recognition, followed by a review of five kinds of EER methods, with an in-depth exploration of several representative approaches.
II-A Problem Formulation
Most current datasets involve recording EEG signals while participants are exposed to specific stimuli, such as videos or music. In our framework, we define a participant in an experiment as a subject. The complete set of EEG data collected from each subject under a particular stimulus condition is referred to as a trial, typically lasting between one to five minutes. Due to the high cost of acquiring EEG datasets, the number of trials within a complete dataset is generally limited. To facilitate subsequent processing tasks, the data within each trial is usually segmented; we denote this segmented data as a sample, which serves as the basic unit for further analysis. Additionally, since the EEG signals of the same subject may fluctuate at different times even when in the same emotional state, some datasets include recordings from various time periods to address the challenges of cross-temporal analysis. We refer to the data collected during a single time period as a session.
Given the definitions of subject, trial, and sample, the goal of emotion recognition is to classify the emotional state of a subject according to EEG samples. These emotional states can be dimensional, such as valence and arousal, or categorical, such as sadness and joy. In an experiment at study, given an EEG dataset $D$, we denote the subjects as $S_{i}(i=1,2,...,n)$, where $n$ refers to the number of subjects within the dataset. For each subject $S_{i}$, the trials they participated in are represented as $T^{(i)}_{j}(j=1,2,...,m)$, where $m$ is the number of trials associated with $S_{i}$. Each trial $T^{(i)}_{j}$ is further divided into samples based on the time dimension.
@@ II-A Problem Formulation
TABLE I: Detailed Information Of Selected Baseline Methods.
Model Baseline frequency Citation count Impact factor of journal Publish time Type
DGCNN[12] 18 1105 11.2 2018 GNN
RGNN[13] 11 545 11.2 2020 GNN
EEGNet[8] 7 2979 4 2018 CNN
DBN[5] 7 1884 2.718 2015 DNN
MS-mda[6] 6 95 5.2 2021 DNN
GCBNet[14] 3 302 11.2 2019 GNN
ACRNN[11] 3 338 11.2 2020 RNN
CDCN[9] 2 142 5 2020 CNN
HSLT[7] 2 79 4.3 2022 Transformer
Tsception[10] 1 136 11.2 2022 CNN
II-B Representative Methods
We based our selection of baseline methods on a comprehensive review of recent high-quality papers in the field of EER from the past few years. After thoroughly examining 60 recent papers, we identified and analyzed 205 distinct baseline methods. Our analysis primarily focused on the frequency with which these methods were used as baselines in the surveyed papers, as this metric best reflects their authority and recognition within the field. Subsequently, we considered additional factors, such as the citation count of each method and the impact factor of the journals (or the rank of conferences) in which they were published, to refine our selection further. Lastly, we ensured a diverse range of methods by taking into account their publication time and methodological variety, including significant methods such as DNN, CNN, RNN, GNN, and Transformer. We filtered out models that were below 10% of those reported in the papers to ensure the validity of the algorithm library comparisons.
Considering the reproducibility of these methods and the limitations of our current resources, we ultimately selected the ten most representative works[8, 9, 10, 11, 12, 13, 14, 5, 6, 7] to include in LibEER. Table I presents detailed information and ranking of these ten methods. These methods represent a well-considered balance of influence, methodological diversity, and publication recency, ensuring comprehensive coverage of the most impactful and up-to-date approaches. In future work, we will continue to maintain and expand this collection by incorporating more recent and diverse methods.
II-B1 CNN-based Methods
Convolutional Neural Networks (CNN) are a type of feedforward neural network characterized by convolutional structure. When applied to EEG signals, CNN treats the EEG data as a 2D representation, where one dimension corresponds to the electrodes and the other dimension represents either temporal or frequency features of the EEG data, as shown in Figure 1 (a).
This paper discusses three key CNN-based methods. EEGNet [8] employs depthwise and separable convolutions across two blocks. The first block features temporal convolution for frequency information and depthwise convolution for spatial filters. The second block uses separable convolution to further refine the features, reducing dimensionality while preserving essential information. Channel-Fused Dense Convolutional Network (CDCN) [9] combines one-dimensional convolution with dense structures to extract temporal and spatial features from EEG signals. Its dense blocks enhance information flow and feature reuse by connecting each layer directly to all preceding layers, effectively capturing temporal dependencies and spatial correlations. TSception [10] is a multi-scale CNN that includes dynamic temporal, asymmetric spatial, and high-level fusion layers. The dynamic temporal layer captures temporal and frequency representations using multi-scale 1D kernels, while the asymmetric spatial layer learns spatial patterns. The high-level fusion layer integrates these features for efficient and accurate emotion classification.
II-B2 RNN-based Methods
Recurrent Neural Networks (RNNs) are designed for sequential data, where directed connections between nodes allow them to maintain a memory of previous inputs. In EEG applications, RNNs extract spatial features by treating channels as sequence nodes and temporal features by treating time slices as sequence nodes, effectively capturing the complex spatiotemporal dynamics in EEG signals (see Figure 1 (b)).
The Attention-based Convolutional RNN (ACRNN) [11] employs channel-wise attention and extended self-attention mechanisms. Channel-wise attention allows the model to focus on relevant channels for emotion recognition, while CNN extracts spatial features from these channels. ACRNN also uses Long Short-Term Memory (LSTM) units to capture and refine temporal features with extended self-attention, assigning importance to different time steps based on their relevance to the recognized emotional state.
II-B3 GNN-based Methods
Graph Neural Network (GNN) is a type of neural network designed to handle data structured as graphs. A graph consists of nodes and edges, which can represent complex relationships and structures. When processing EEG data, the EEG channels are typically treated as nodes in a graph, as shown in Figure 1 (c). Since EEG data does not inherently contain information about the edges or the adjacency matrix, using graph neural networks to process EEG data often involves leveraging prior knowledge about the relationships between channels or using learnable adjacency matrices as the graph’s adjacency matrix.
II-B4 DNN-based Methods
Deep Neural Network (DNN) is a feedforward artificial network consisting of fully connected neurons with a nonlinear activation function, including the input layer, multiple hidden layers, and the output layer. In the field of EER, DNN processes EEG data by transforming and mapping it into the feature space, enabling the extraction and utilization of complex patterns within the data, illustrated in Figure 1 (d).
II-B5 Transform-based Methods
Transformer uses a self-attention mechanism to capture dependencies between elements in a sequence, effectively handling long-range dependencies. As shown in Figure 1 (e), the EEG data are typically segmented into multiple patches on channels, regions, or temporal segments when using transformers to process EEG data. These patches are then fed into the transformer model for further analysis.
TABLE II: Summary Of Our Proposed Benchmark.
Module Key Information
Datasets SEED / SEED-IV / DEAP / MAHNOB-HCI
Experimental Task Subject-dependent / Cross-subject
Data Splitting Train : Val : Test = 0.6 : 0.2 : 0.2
Evaluation Methods Accuracy / F1-score
@@ II-B5 Transform-based Methods
III Benchmark Building
A comprehensive, rational, and transparent benchmark is essential for the equitable comparison of research methodologies within the field, a need currently unmet in the EER domain. In this section, we will elaborate on the benchmark we have devised, encompassing datasets, data preprocessing, data partitioning, and evaluation methods. For convenience, we summarize the key information of the EER benchmark in Table II.
III-A Datasets
Research in the EER field exhibits variability in dataset selection and often lacks clarity regarding the specific data segments utilized. This study evaluates the performance of popular models using four commonly employed datasets, which will be introduced along with the relevant data segments, as summarized in Table III.
III-A1 SEED

### v2

@@ Abstract
EEG-based emotion recognition (EER) has gained significant attention due to its potential for understanding and analyzing human emotions. While recent advancements in deep learning techniques have substantially improved EER, the field lacks a convincing benchmark and comprehensive open-source libraries. This absence complicates fair comparisons between models and creates reproducibility challenges for practitioners, which collectively hinder progress. To address these issues, we introduce LibEER, a comprehensive benchmark and algorithm library designed to facilitate fair comparisons in EER. LibEER carefully selects popular and powerful baselines, harmonizes key implementation details across methods, and provides a standardized codebase in PyTorch. By offering a consistent evaluation framework with standardized experimental settings, LibEER enables unbiased assessments of over ten representative deep learning models for EER across the four most widely used datasets. Additionally, we conduct a thorough, reproducible comparison of model performance and efficiency, providing valuable insights to guide researchers in the selection and design of EER models. Moreover, we make observations and in-depth analysis on the experiment results and identify current challenges in this community. We hope that our work will not only lower entry barriers for newcomers to EEG-based emotion recognition but also contribute to the standardization of research in this domain, fostering steady development. The library and source code are publicly available at https://github.com/XJTU-EEG/LibEER.
@@ I Introduction
Emotion represents a spectrum of subjective cognitive experiences that forms the core framework for interpersonal relationships and serves as a foundation for evaluating human behaviors [1]. Due to its non-invasive nature, high temporal resolution, portability, and cost-effectiveness, electroencephalography (EEG) has become a prominent physiological signal in emotion analysis [30, 31]. As a result, EEG-based emotion recognition (EER) has attracted significant research interest in recent years, with applications spanning diverse fields such as healthcare [32, 41], advertising [33], and education [34].
Traditional approaches to EER primarily rely on hand-crafted features, such as power spectral density (PSD)[40] and differential entropy (DE)[16]. Recently, with advancements in automatic feature engineering, deep learning models have emerged as a promising research area in EER and achieved dominant performance [17, 19, 2]. These models generally target to extract temporal features within each EEG channel, spatial features across different EEG channels, and spectral features in the frequency domain.
Despite significant advancements in deep learning model design, the practical implementation of these approaches is often overlooked, making it difficult for newcomers and researchers to keep up with emerging studies. Therefore, this work focuses on the implementation aspects of EER. The specific issues we aim to address are as follows:
  • I1 (Lack of benchmarking in EER): A common practice in many model design studies is to compare new approaches with state-of-the-art (SOTA) models. However, such comparisons are often unreliable due to the absence of comprehensive benchmarking within the EER community. Studies frequently use varying preprocessing techniques, such as artifact removal, and different preprocessing settings, such as varying train/test split ratios. Some studies also adopt questionable practices, such as trial-dependent configurations [35, 11], which can introduce data leakage risks. Furthermore, many studies provide ambiguous experimental settings [7, 36, 37], which hinder reproducibility. To address these issues, standardizing preprocessing techniques and experimental settings is essential for enabling fair comparisons and identifying genuine progress in the field.
  • I2 (Lack of open-source libraries for deep learning-based EER): Implementation details, such as training hyperparameters, play a crucial role in model performance. While it is impractical and unnecessary to include every detail in an academic paper, making implementation code publicly available is vital for enabling the reproduction and further development of SOTA research. However, even when code is provided, it is often tied to different platforms, such as PyTorch [20], TensorFlow [21], or MATLAB [27], posing challenges to researchers. Therefore, there is a pressing need for an open-source library based on a standardized platform that provides comprehensive implementation details while achieving performance comparable to the original study.
  • I3 (Lack of in-depth analysis of SOTA EER research): While deep learning has dominated EER research, many survey papers  have been published summarizing progress from various perspectives. However, these surveys typically report findings without hands-on experience in reproducing results, which may lead to misdirected future research. This lack of practical insight can obscure real challenges faced in the field. Therefore, an in-depth analysis of SOTA EER research based on reproduction results is essential for uncovering the true challenges and advancing the field.
In response to these challenges, we propose LibEER (Library for EEG-based Emotion Recognition) and conduct an analysis of SOTA EER research based on reproduction results with LibEER. To establish a comprehensive benchmark in EER (addressing issue I1), we standardize datasets, evaluation metrics, and experimental settings. Specifically, we review recent SOTA EER studies to identify the most commonly used datasets, SEED [23], SEED-IV [25], DEAP [22], and MAHNOB-HCI [24], and evaluate the corresponding metrics. Our benchmark includes both subject-dependent and subject-independent settings, along with a variety of preprocessing methods. In contrast to the simplified train/test split commonly used in the literature, we adopt a more rigorous train/validation/test split to ensure a more reliable performance evaluation. To address issue I2, we develop the LibEER by systematically selecting and implementing ten representative deep learning-based EER methods, all within a unified training pipeline in PyTorch. LibEER offers a wide range of EEG signal preprocessing functions, evaluation protocols, and experimental scenarios. To enhance its flexibility, we provide user-friendly interfaces that allow researchers to easily customize experimental configurations. Users can set hyperparameters, choose preprocessing methods, and modify backbone models to suit their specific needs. Based on the reproduction results from LibEER, we conduct a series of comprehensive comparison experiments and identify four key observations (addressing issue I3). Through these experiments, we highlight both the real limitations and the progress made in current studies.
@@
It is worth noting that recent research [3] has made important contributions by providing a toolbox for EER. However, our work differs significantly and builds upon these contributions in several key areas. First, we systematically analyze the challenges associated with making fair comparisons in EER tasks, a discussion notably absent in [3]. Second, while the existing benchmark [3] includes a variety of methods, many of them are not specifically tailored for EER tasks. In contrast, we have selected the ten most representative deep learning-based EER methods, based on their impact in this field, making our benchmark better reflect the true SOTA. Finally, the existing benchmark [3] predominantly focuses on research that has already made its code publicly available. However, many influential EER studies do not provide open-source code, which has led to the exclusion of some significant models in their work. In response, we have carefully reproduced all representative studies, including those without publicly available code and rigorously validated the results to ensure their accuracy. We believe that our benchmark and library offer fresh perspectives that will significantly advance the field.
In summary, our main contributions are as follows:
@@
  • We have pioneered the establishment of the first benchmark specifically designed to enable fair comparisons in EEG-based emotion recognition (EER). This benchmark not only reveals the true progress of the EER task but also provides a convenient reference for conducting comparative experiments.
@@
  • We have introduced LibEER as an open-source project on GitHub, offering flexible customization of experimental schemes through user-friendly interfaces. LibEER supports fair and comprehensive evaluation of over ten popular models across four datasets. Serving both as a toolbox and a platform, it assists practitioners in efficiently utilizing and reproducing EER methods.
@@
  • We provide thorough evaluations across various experimental settings, carefully controlling implementation details. These experimental results offer valuable insights into the advancements made and assist researchers in selecting or designing EER models.
The paper is organized as follows. Section II provides the background of the EER task. Section III presents the design of the benchmark. Section IV details the library. Section V covers experimental results and in-depth analysis. Section VI identifies future work, and Section VII concludes the whole work.
Figure 1: Data collection protocol of EER.
II Background of EER
In this section, we present the general background of EEG, including data collection protocol and formal problem definition.
II-A Data Collection Protocol
The general protocol for EER data collection is illustrated in Fig. 1. Each data collection begins with a hint of start shown to the participant. Afterwards, various stimuli, such as images, audio, or video, are presented to evoke specific emotions in the participant, such as happiness, sadness, or fear. After each stimulus, the participant is asked to self-assess their emotional state, providing a subjective emotion score based on their experience during the stimulus. This process is repeated a set number of times with different stimuli to collect enough data.
In this context, the participant exposed to the stimuli, from whom data is subsequently collected, is commonly referred to as a subject. A continuous segment of EEG data collected from a subject under a specific stimulus condition is called a trial, typically lasting between one and five minutes. Between trials, the subject takes a rest. A series of trials is termed a session. Since EEG data from the same subject, even in the same emotional state, may vary significantly over time, datasets generally include several sessions, with an interval of more than 24 hours between consecutive ones. Given the high cost of EEG data acquisition, a trial is often divided into segments. Two consecutive segments within a trial can be overlapping or non-overlapping. Each segment is called a sample, which serves as the fundamental unit for EER. Notably, within a single trial, all samples should be designated for either training or test. A trial-dependent setting, where some samples within a trial are used for training and others for testing, is impractical for real-world applications and should be considered an invalid setup.
II-B Problem Formulation
Given an EER dataset $D$, a sample is denoted as $x_{i,j,k,q}\in\mathbb{R}^{c\times d}$, where $i$, $j$, $k$, and $q$ refer to the $i$-th subject, $j$-th session, $k$-th trial, and $q$-th sample within the trial, respectively. Here, $c$ represents the number of electrode channels, and $d$ denotes the feature dimension of each sample. The goal of EER is to classify a sample into a designated emotional state, which may be represented dimensionally (e.g., valence, arousal) or categorically (e.g., sadness, joy), via the following mapping:
$\mathcal{F}:x\rightarrow\boldsymbol{y},$ (1)
where $\boldsymbol{y}$ denotes the emotional state and $\mathcal{F}$ is a mapping function. SOTA map functions are usually neural networks trained on a training set $D_{\text{train}}\subset D$, and evaluated on a test set $D_{\text{test}}\subset D$. Additionally, a validation set $D_{\text{valid}}\subset D$ is often employed to select the best model. These subsets are required to be mutually exclusive, that is $D_{\text{train}}\cap D_{\text{test}}\cap D_{\text{valid}}=\emptyset$.
TABLE I: Summary of our proposed benchmark.
Module Key Information
Data Splitting Train : Val : Test = 0.6 : 0.2 : 0.2
Experimental Task Subject-dependent / Cross-subject
Datasets SEED / SEED-IV / DEAP / MAHNOB-HCI
Evaluation Metrics Accuracy / F1-score
TABLE II: Detailed information of selected baseline methods.
Model Baseline frequency Citation count Publication Publish year Type
Tsception[10] 2 136 TAFFC 2022 CNN
CDCN[9] 2 142 TCDS 2020 CNN
EEGNet[8] 7 2979 JNE 2018 CNN
R2G-STNN[14] 8 235 TAFFC 2022 RNN
BiDANN[36] 4 250 TAFFC 2021 RNN
ACRNN[11] 3 338 TAFFC 2020 RNN
MS-mda[6] 6 95 Front. Neurosci. 2021 DNN
DBN[5] 7 1884 TAMD 2015 DNN
RGNN[13] 11 545 TAFFC 2020 GNN
GCBNet[15] 3 302 TAFFC 2019 GNN
DGCNN[12] 18 1105 TAFFC 2018 GNN
HSLT[7] 2 79 IEEE SENS J 2022 Transformer
III Benchmark Building
A comprehensive, well-structured, and transparent benchmark is essential for fair comparisons of research methodologies in the field—a need that remains unmet in the EER domain. In this section, we present our benchmark, covering baseline selection, preprocessing, experimental setups, datasets, and evaluation protocols. For ease of reference, TABLE I provides a summary of the key information in our EER benchmark.
III-A Baseline Selection
We select representative models for the benchmark. Specifically, we set up the following criteria:
  • We first collected a total of 60 high-quality papers published in the past year in the field of EEG-based Emotion Recognition (EER) from reputable journals and conferences (such as TAFFC, JBHI, and ICASSP).
  • We then compiled a list of 205 comparative methods cited across these 60 papers.
  • After that, we ranked these 205 methods based on their frequency as comparison methods in the 60 papers and selected the top 20 for replication.
  • These 20 methods, based on model architecture, are divided into five categories: CNN, RNN, GNN, DNN, and Transformer. Finally, we selected models with replication results within a 10% error margin compared to the original reported results, while ensuring that each category includes at least one representative model.
Based on these steps, we ultimately identified the twelve most representative works [8, 9, 10, 11, 12, 13, 15, 5, 6, 7, 36, 14] to include in LibEER. TABLE II presents detailed information on these twelve methods. In future work, we will continue to maintain and expand this collection by incorporating more recent and diverse methods. The following subsections will offer detailed descriptions of the five categories, as well as an examination of each selected method.
III-A1 CNN-based Methods
Convolutional Neural Networks (CNNs) are a type of feedforward neural network characterized by their convolutional operations. When applied to EEG signals, CNNs usually treat EEG data as a two-dimensional (2D) representation: one dimension corresponds to the channels, while the other dimension represents either the temporal or spectral features of the EEG data. This structure is illustrated in Fig. 2 (a), which could handle temporal/spectral and spatial information simultaneously. On the other hand, some work relies on 1D CNN to extract temporal and spatial features separately.
LibEER includes three representative CNN-based methods: EEGNet [8], CDCN [9], and TSception [10]. Although not initially proposed for EER, EEGNet is included as a representative 2D CNN-based model due to its significant influence, with over 3000 citations. Both CDCN and TSception employ 1D CNNs for spatial and temporal feature extraction. TSception, a more advanced approach, incorporates multi-scale 1D CNNs for temporal features and employs hemisphere kernels for spatial features.
III-A2 RNN-based Methods
Recurrent Neural Networks (RNNs) are inherently designed for sequential data, making them well-suited for extracting temporal features from EEG signals. For spatial feature extraction, RNNs treat channels as sequence nodes, as illustrated in Fig. 2 (b). In practice, Long Short-Term Memory (LSTM) is a widely used variant of RNNs.
Our benchmark includes three RNN-based methods: ACRNN [11], BiDANN [36], and R2G-STNN [14]. Both ACRNN and BiDANN use LSTM for temporal feature extraction, while R2G-STNN applies bidirectional LSTM to capture spatial features. An attention mechanism is often integrated with LSTM to enhance feature extraction; for instance, ACRNN uses channel-wise attention for spatial features and extended self-attention with LSTM for temporal feature extraction.
III-A3 GNN-based Methods
Graph Neural Network (GNN) is a type of neural network designed to handle data structured as graphs. A graph consists of nodes and edges, which can represent complex relationships and structures. When processing EEG data, the EEG channels are typically treated as nodes in a graph, as shown in Fig. 2 (c). Since EEG data does not inherently contain information about the edges or the adjacency matrix, using graph neural networks to process EEG data often involves leveraging prior knowledge about the relationships between channels or using learnable adjacency matrices as the graph’s adjacency matrix.
III-A4 DNN-based Methods
Deep Neural Network (DNN) is a feedforward artificial network consisting of fully connected neurons with a nonlinear activation function, including the input layer, multiple hidden layers, and the output layer. In the field of EER, DNN processes EEG data by transforming and mapping it into the feature space, enabling the extraction and utilization of complex patterns within the data, illustrated in Fig. 2 (d).
III-A5 Transform-based Methods
Transformer uses a self-attention mechanism to capture dependencies between elements in a sequence, effectively handling long-range dependencies. As shown in Fig. 2 (e), the EEG data are typically segmented into multiple patches on channels, regions, or temporal segments when using transformers to process EEG data. These patches are then fed into the transformer model for further analysis.
III-B Data preprocessing and Splitting
III-B1 preprocessing
III-B2 Splitting
III-C Experimental Tasks
The subject-independent task lies between the subject-dependent and cross-subject tasks, with the latter two being more aligned with real-world applications. The cross-session task necessitates datasets with multiple sessions, which many datasets lack. Therefore, our benchmark focuses on subject-dependent and cross-subject tasks, as these are the most commonly utilized and hold significant practical relevance. Additionally, it is important to note that we treat data from different sessions of the same subject as if they originate from distinct subjects. We provide specific definitions for the subject-dependent and cross-subject tasks below.
@@ III-C Experimental Tasks
Figure 3: The framework and data pipeline of LibEER
III-D Datasets
Research in the EER field exhibits variability in dataset selection and often lacks clarity regarding the specific data segments utilized. This study evaluates the performance of popular models using four datasets, which are selected by comprehensively considering factors such as their frequency of use, data scale, and data quality. They will be introduced along with the relevant data segments, as summarized in Table III.
III-D1 SEED